# Robust $\mathcal{H}_{\infty}$ control under stochastic requirements: minimizing conditional value-at-risk instead of worst-case performance

### v2

@@ V-A Benchmark
The uncertain parameters represent mechanical properties. Once normalized, they all follow a Gaussian distribution of mean 0 and standard deviation $1/3$, truncated at 3 $\sigma$, except for one parameter (representing an angle) which follows a uniform distribution between -1 and 1. Additionally, the vectors $\delta$ that do not satisfy a nonlinear constraint $c(\delta)\leq 0$, related to some mechanical constraint detailed in [15], must be removed from the robustness analysis. In the deterministic setting, this constraint can be addressed in the framework of constrained optimization as in [15]. In the $\beta$-CVAR formulation, when performing Monte-Carlo sampling, samples that do not satisfy the constraint are simply discarded; all following results ($\beta$-VAR, $\beta$-CVAR, mean, distribution) refer to the resulting conditional distribution.
@@
V-B Formulation as a deterministic control problem
@@
$\mathcal{D}_{\delta}=\{\delta\in[-1,1]^{m}\text{ s.t. }c(\delta)\leq 0\}$
@@
and is solved with the method of [15]. Let us note $\hat{k}_{det}$ the decision parameters describing the optimal controller. The resulting densities for $L_{un}(s,\hat{k}_{det},\delta)$ (Req. 1) and $L_{us}(s,\hat{k}_{det},\delta)$ (Req. 2) are presented in Fig. 4 using 10000 samples. It can be verified that the worst case for Req. 2, computed with the deterministic method of [15], is indeed inferior to 1, as imposed. It is also verified for $L_{er}(s,\hat{k}_{det},\delta)$ (Req. 3), not represented for brevity (but the performance metrics are all summarized in Table I). Nominal performance, i.e. when $\delta=0$, is also indicated for completeness.
@@
V-C Formulation as a stochastic control problem
Based on the notations of Definition 1, we fix $\beta=95$% and seek to solve the stochastic problem:
@@ V-C Formulation as a stochastic control problem
We solved this problem with Algorithm 1, using $\hat{k}_{\textrm{det}}$ as initial controller (step 1). Step 2 was performed with Matlab’s function fmincon using the sequential quadratic programming (SQP) algorithm, even though convergence to an optimum is not formally guaranteed due to the nonsmoothness of the SAA. 2500 samples were used initially, and this number was increased to 10000 following step 4. With these 10000 samples, the 99% interval of confidence of Eq. (13) was up to $0.0008$ among the 3 requirements, which was considered sufficient for validation. Stability over the whole domain was also verified with [15]. We note $\hat{k}_{sto}$ the decision parameters describing the optimal controller.
To compare with the results of the deterministic formulation of Section V-B, we also computed the worst case with the method of [15], although this computation was not performed during the optimization itself. The results are presented in Fig. 5 and summarized in Table I. In particular:
1) Fig. 5(b): The $\beta$-CVAR of $L_{us}(s,\hat{k}_{sto},\delta)$ (Req. 2) is 0.98, which is inferior to 1 as imposed. A worst-case of 2.05 (again found with the deterministic method of [15]) is tolerated, whereas it would have been rejected with the traditional formulation. Nevertheless, this worst case is very rare. To illustrate this statement, in addition to the risk measures already presented (VAR, CVAR), we mention that the worst-case performance found among the 1000010000 samples used to draw the histogram is 1.58, and already covers at least 99.9% of the distribution with a confidence of 99.99%. [^2] For a random function $f(\delta)$, noting $\delta^{*}$ the sample realizing the largest function evaluation, [27] shows that $\mathrm{P}(\mathrm{P}(f(\delta)>f(\delta^{*})\leq\epsilon)\geq 1-\gamma$ is guaranteed when NN verifies: $N>\frac{\textrm{ln}(\gamma)}{\textrm{ln}(1-\epsilon)}\;.$
2) Fig. 5(a): As a result of relaxing Req. 2 (from worst case to CVAR criterion), the Req. 1 is considerably improved: all metrics were reduced by $\sim$30%.
@@
TABLE I: Performance metrics for the deterministic (det.) and stochastic (sto.) approaches.
@@
VI Conclusion
To reduce the conservatism of conventional robust $\mathcal{H}_{\infty}$ control, which minimizes a worst-case performance that may be very unlikely, we introduced an alternative stochastic approach based on the conditional value at risk. Its capacity to significantly improve the performance of uncertain control systems, at the expense of tolerating rare degrading worst-case configurations, was illustrated on a mechanical system.
ACKNOWLEDGMENT
The authors acknowledge the use of an AI language model (ChatGPT: GPT-5.2) to assist in improving the wording and reformulation of selected sentences.

### v3

@@ 5.1 Benchmark
The uncertain parameters represent mechanical properties. Once normalized, they all follow a Gaussian distribution of mean 0 and standard deviation $1/3$, truncated at 3 $\sigma$, except for one parameter (representing an angle) which follows a uniform distribution between -1 and 1. Additionally, the vectors $\delta$ that do not satisfy a nonlinear constraint $c(\delta)\leq 0$, related to some mechanical constraint detailed in [16], must be removed from the robustness analysis. In the deterministic setting, this constraint can be addressed in the framework of constrained optimization as in [16]. In the $\beta$-CVAR formulation, when performing Monte Carlo sampling, samples that do not satisfy the constraint are simply discarded; all following results ($\beta$-VAR, $\beta$-CVAR, mean, distribution) refer to the resulting conditional distribution.
@@
5.2 Formulation as a deterministic control problem
@@
$\mathcal{D}_{\delta}=\{\delta\in[-1,1]^{m}\text{ such that }c(\delta)\leq 0\}$
@@
and is solved with the method of [16]. Let us note $\hat{k}_{det}$ the decision parameters describing the optimal controller. The resulting densities for $L_{un}(s,\hat{k}_{det},\delta)$ (Req. 1, Fig 4(a)) and $L_{us}(s,\hat{k}_{det},\delta)$ (Req. 2, Fig 4(b)) are presented in Fig. 4 using 10000 samples. It can be verified that the worst case for Req. 2, computed with the deterministic method of [16], is indeed less than 1, as imposed. The mean, $\beta$-VaR and $\beta$-CVaR ($\beta$ = 95%) are reported even though they were not used when solving the optimization problem. Nominal performance, i.e. when $\delta=0$, is also indicated for completeness. All performance metrics are summarized in Table 1, including Req. 3, which is not represented in the figures.
@@
5.3 Formulation as a stochastic control problem
Based on the notations of Definition 3.1, we fix $\beta=95$% and seek to solve the stochastic problem:
@@ 5.3 Formulation as a stochastic control problem
We solved this problem with Algorithm 1, using $\hat{k}_{\textrm{det}}$ as initial controller (step 1). Step 2 was performed with MATLAB’s function fmincon [12] using the sequential quadratic programming (SQP) algorithm, even though convergence to an optimum is not formally guaranteed due to the nonsmoothness of the SAA. The constraint on the spectral abscissa was not explicitly taken into account when solving Step 2, but stability over the whole domain was verified a posteriori in Step 3 using the method of [16] (note that Req. 2, by imposing a constraint on the CVAR of the stability margins, helps enforce stability). 2500 samples were used initially in the SAAs, and this number was increased to 10000 following step 4. With these 10000 samples, the 99% confidence interval of Eq. (13) was at most $0.0008$ among the 3 requirements, which was considered sufficient for validation. We note $\hat{k}_{sto}$ the decision parameters describing the optimal controller.
To compare with the results of the deterministic formulation of Section 5.2, we also computed the worst case with the method of [16], although this computation was not performed during the optimization itself. The results are presented in Fig. 5 and summarized in Table 1. In particular:
1) Fig. 5(b): The $\beta$-CVAR of $L_{us}(s,\hat{k}_{sto},\delta)$ (Req. 2) is 0.98, which is less than 1 as imposed. A worst-case of 2.05 (again found with the deterministic method of [16]) is tolerated, whereas it would have been rejected with the traditional formulation. Nevertheless, this worst case is very rare. To illustrate this statement, we also mention that among the 10,000 samples used to construct the histogram, the maximum observed value was 1.58. This implies, with 99.99% confidence, that at least 99.9% of the underlying distribution lies below 1.58.222For a random function $f(\delta)$, noting $\delta^{*}$ the sample realizing the largest function evaluation, [28] shows that $\mathrm{P}(\mathrm{P}(f(\delta)>f(\delta^{*}))\leq\epsilon)\geq 1-\gamma$ is guaranteed when $N$ verifies: $N>\frac{\textrm{ln}(\gamma)}{\textrm{ln}(1-\epsilon)}\;.$
2) Fig. 5(a): As a result of relaxing Req. 2 (from worst case to CVAR criterion), the Req. 1 is considerably improved: all metrics were reduced by $\sim$30% with respect to Section 5.2.
@@
Table 1: Performance metrics for the deterministic (det.) and stochastic (sto.) approaches.
@@
6 Conclusion
To reduce the conservatism of conventional robust $\mathcal{H}_{\infty}$ control, which minimizes a worst-case performance that may be very unlikely, we introduced an alternative stochastic approach based on the conditional value at risk. Its capacity to significantly improve the performance of uncertain control systems, at the expense of tolerating rare degrading worst-case configurations, was illustrated on a mechanical system. Finally, we presented a first step toward a synthesis method and highlighted several remaining open challenges.